Trace the curve 3 2 y = x −1 , and state all the properties you use to trace it.
1
Expert's answer
2020-03-10T13:22:49-0400

32y=x-1

y=(x-1)/32

This is a graph of a function of the straight line in the form y = kx + b

It is defined on the set of real numbers and monotonically increases.


The straight line goes through the points (1;0), (0; -0.03125).